OpenStudy (anonymous):

what is log (base sqrt 3) 729=x

OpenStudy (amistre64):

sqrt(3)^x = 729 is what it means .... 3^(1/2)^x = 729 3^(x/2) = 729 so... what does 3^a = 729 get us :)

OpenStudy (amistre64):

doing these by hand can be tedious, the quick way is just to change of base it and calculte

OpenStudy (amistre64):

x = 12 :)

OpenStudy (amistre64):

change of base goes: logB(A) = x ln(A) ---- = x ln(B)

OpenStudy (anonymous):

ok i worked thru it with u and got it...the whole sqrt thing just confuses me sometime lol